Factor the trinomial (can you please show & tell me how to do this)
5tsquared -22t-15

Answer by kkasko(46)

Factor the trinomial (can you please show & tell me how to do this)
The first step is to find _x_= 5x-15=-75
_+_=-22

5t^2 -22t-15. -25+3=-22. -25x3=-75
Now put the -25t+3t into the equation where -22t was so.
5t^2-25t+3t-15
Now break the trinomial into two factors by 5x_=-25 and after the bracket put a 3 and then another bracket( t and 3x_=-15
5t(t-5)3(t-5)
(5t+3)(t-5)
These are the factors to the trinomial